# Measurement Error and Counterfactuals in Quantitative Trade and Spatial Models

### v4

@@ 3.2 Introducing Measurement Error
we can use Bayes’ rule to find the posterior distribution of the true data given the noisy data,
@@
This posterior distribution then allows us to generate draws from our posterior distribution for the true data given the noisy data. [^5] Note that the measurement error distribution does not have to be mean zero, so also allows for measurement error bias. Nevertheless, even mean zero measurement error can result in bias in the counterfactual prediction of interest. This is automatically taken into account by the Bayesian approach when quantifying uncertainty. Furthermore, the individual measurement error distributions can be arbitrarily correlated in this general setup.
The Bayesian approach allows researchers to incorporate economic knowledge through the prior. For example when considering measurement error in non-negative flows between locations one can fit a prior centered on a gravity model, which I will do in Section 4.
3.2.1 Running Example: Armington Model (Continued)
For the Armington model, I will assume that there is measurement error in trade flows $\left\{F_{ij}\right\}$. This is plausible because in Linsi, Burgoon, and Mügge (2023) it is shown that there are so-called mirror discrepancies in bilateral trade flows between almost all countries. This means that, for instance, while the value that Germany reports it imported from France and the value that France reports it exported to Germany should be the same, in practice they are often different.
Hence, instead of the true trade flows we observe noisy trade flows $\left\{\tilde{F}_{ij}\right\}$, which in turn lead to noisy counterfactual predictions $\tilde{\gamma}_{q}$ for $q\in\left\{\mathrm{CAF},\mathrm{NLD},\mathrm{SWE},\mathrm{USA}\right\}$. If we specify a prior $\pi\left(\left\{F_{ij}\right\}\right)$ and a measurement error model $\pi\left(\left\{\tilde{F}_{ij}\right\}|\left\{F_{ij}\right\}\right)$, we can use Bayes’ rule to find the posterior $\pi^{ME}\left(\left\{F_{ij}\right\}|\left\{\tilde{F}_{ij}\right\}\right)$.
The default approach presented later in Section 4 can be applied to this setting, so we can use the provided toolkit to obtain draws from $\pi^{ME}\left(\left\{F_{ij}\right\}|\left\{\tilde{F}_{ij}\right\}\right)$. Fixing the structural parameter at its point estimate, we can see the impact of measurement error in Table 2 and Figure 2. For Sweden, measurement error does not have that much of an impact compared to estimation error. For the Central African Republic the variance due to measurement error is comparable to the variance due to estimation error, but accounting for measurement error shifts the posterior towards zero relative to the posterior that only accounts for estimation error. This indicates that measurement error caused a bias here. For the Netherlands and the United States, measurement error causes much more uncertainty, as the variance of the posterior accounting for measurement error is much larger than the variance of the posterior accounting for estimation error. Furthermore, for the Netherlands there is a considerable bias correction. These plots illustrate that the proposed approach automatically incorporates bias that is caused by measurement error.
Point estimate Only est. error Only meas. error
$\gamma_{\mathrm{CAF}}$0.72 [-0.95, -0.13] [-0.46, -0.05]
$\gamma_{\mathrm{NLD}}$5.47 [-5.50, -5.37] [-7.91, -6.28]
$\gamma_{\mathrm{SWE}}$3.77 [-4.14, -3.54] [-4.24, -3.46]
$\gamma_{\mathrm{USA}}$1.09 [-1.09, -1.03] [-1.41, -0.36]
Table 2: Uncertainty quantification for the Armington model, considering estimation error and measurement error separately.
Figure 2: Uncertainty quantification for the Armington model, considering estimation error and measurement error separately.
3.2.2 Relation to Measurement Error Literature
The literature on measurement error in nonlinear models is extensive, as reviewed in Hu (2015) and Schennach (2016), and the most closely related strand of measurement error literature is that on nonseparable error models (Matzkin, 2003; Chesher, 2003; Hoderlein and Mammen, 2007; Matzkin, 2008; Hu and Schennach, 2008; Schennach, White, and Chalak, 2012; Song, Schennach, and White, 2015). However, these results do not apply to my setting. The key distinguishing feature of the setting in this paper is that the object of interest $\gamma$ directly depends on the correctly measured data, because the equality in Assumption 1 is an exact statement. In contrast, in conventional measurement error settings the object of interest is a function of the correctly measured distribution of the data, $\mathcal{P}_{D}$, rather than the actual realized observations, $D$. This leads to the key distinction in Equation (2).
This difference is important because in my setting, it would not suffice to be able to perfectly estimate the distribution $\mathcal{P}_{D}$. For example in the running example, to answer counterfactual questions we need the realized trade flows $\left\{F_{ij}\right\}$, rather than the trade flow distribution from which they are drawn. In contrast, in a conventional measurement error setting knowing this distribution would suffice, because the estimands are functionals of the correctly measured distribution of the data. By virtue of that, we need to account for uncertainty about the observations themselves rather than their distribution.
3.3 Quantifying Uncertainty about $\gamma$
The object of interest is a function of the true data and the structural parameter. From the discussion in the previous sections, it follows that we must consider estimation error, the direct effect of mismeasurement, and the indirect effect of mismeasurement through the estimation procedure. Our goal is to quantify uncertainty about $\gamma$ when we observe $\tilde{D}$ by accounting for these various sources of uncertainty.
Recall that we have obtained two different posteriors. The first one is the posterior distribution of $\gamma$ given the true data, $\pi^{EE}\left(\gamma|D\right)$, which incorporates estimation error. The second one is the posterior of the true data given the noisy data, $\pi^{ME}\left(D|\tilde{D}\right)$, which incorporates measurement error. We can combine these two posteriors in different ways to quantify uncertainty about $\gamma$. Here, I will use $\mathbb{E}_{\pi}$ and $Pr_{\pi}$ to denote the expectation and probability under a posterior $\pi$, respectively.
The first approach aims to find an interval $\mathcal{C}^{1}$ to which, in posterior expectation over $D$, the posterior $\pi^{EE}\left(\gamma|D\right)$ assigns probability $1-\alpha$:
$\mathbb{E}_{\pi^{ME}}\left[Pr_{\pi^{EE}}\left\{\gamma\in\mathcal{C}^{1}|D\right\}|\tilde{D}\right]\geq 1-\alpha.$
@@ 3.3 Quantifying Uncertainty about $\gamma$
In practice, given $\tilde{D}$ one would generate draws from $\pi^{ME}\left(D|\tilde{D}\right)$, and for each of these draws obtain a corresponding draw from $\pi^{EE}\left(\gamma|D\right)$. [^6] If an estimator from another study is used, then $\pi^{EE}\left(\theta|D\right)\approx\mathcal{N}\left(\tilde{\theta},\tilde{\Sigma}\right)$. In that case, we can sample $\theta_{b}$ and $D_{b}$ separately, which makes the algorithm much faster. Then, one would report the $\alpha/2$ and $21-\alpha/$ 2 quantiles of this second set of draws. [^7] Note that counterfactual predictions are typically derived as functions of the full system of counterfactual equilibrium variables. Thus, whether the researcher is ultimately interested in a scalar outcome, a relative comparison, or a global average, the mechanics of uncertainty quantification—drawing from the posterior over the true data and parameters and solving for equilibrium—remain the same. This is summarized in Algorithm 1.
@@
    Input: noisy data $\tilde{D}$, number of bootstrap draws $B$, coverage level $1-\alpha$ (choose $B$ and $\alpha$ such that $\alpha/2\cdot B\in\mathbb{N}$).
@@
    For $b=1,...,B$,
@@
      Sample $\theta_{b}\sim\pi^{EE}\left(\theta|D_{b}\right)$.
    3. (c)
      Compute $\gamma_{b}=g\left(D_{b},\theta_{b}\right).$
  3. 3.
@@
The second approach is more conservative. Suppose we again obtain draws from the posterior $\pi^{ME}\left(D|\tilde{D}\right)$ and for each draw use $\pi^{EE}\left(\gamma|D\right)$ to compute an interval that covers $\gamma$ with probability $1-\alpha$. The second approach then aims to find an interval $\mathcal{C}^{2}$ that covers these $100\left(1-\alpha\right)\%$intervals with probability $1-\alpha$:
$Pr_{\pi^{ME}}\left\{Pr_{\pi^{EE}}\left\{\gamma\in\mathcal{C}^{2}|D\right\}\geq 1-\alpha|\tilde{D}\right\}\geq 1-\alpha.$
In practice, one can generate $100\left(1-\alpha\right)\%$intervals around $\gamma$ for many draws from $\pi^{ME}\left(D|\tilde{D}\right)$, and then report the $\alpha/2$ quantile of the set of lower bounds and the $21-\alpha/$ 2 quantile of the set of upper bounds.
Going forward, I will focus on the less conservative interval $\mathcal{C}^{1}$, because it will turn out that in applications the interval $\mathcal{C}^{2}$ will yield extremely wide intervals for many cases. In Appendix G.3 I compute the interval $\mathcal{C}^{2}$ for one of my applications. [^8] Note that one could in principle use a single prior $\pi$ on the underlying data generating process to handle both estimation error and measurement error. I instead combine two simple priors to separately handle estimation error and measurement error, since this leads to highly tractable procedures, albeit at the cost of complicating the Bayesian interpretation of resulting intervals.
3.3.1 Running Example: Armington Model (Continued)
Consider again the Armington model with noisily measured bilateral trade flows. Table 3 and Figure 3 display the resulting intervals and posterior distributions for the objects of interests, respectively. The posteriors that account for both estimation and measurement error are compositions of the two previously plotted posterior distributions. In addition, Figure 4 presents the posterior distributions for the trade elasticity. We observe that measurement error induces attenuation bias.
Point estimate Only est. error Only meas. error Est. and meas. error
$\gamma_{\mathrm{CAF}}$0.72 [-0.95, -0.13] [-0.46, -0.05] [-0.47, -0.03]
$\gamma_{\mathrm{NLD}}$5.47 [-5.50, -5.37] [-7.91, -6.28] [-8.08, -6.24]
$\gamma_{\mathrm{SWE}}$3.77 [-4.14, -3.54] [-4.24, -3.46] [-4.32, -3.45]
$\gamma_{\mathrm{USA}}$1.09 [-1.09, -1.03] [-1.41, -0.36] [-1.44, -0.31]
Table 3: Uncertainty quantification for the Armington model, considering estimation error and measurement error simultaneously.
Figure 3: Uncertainty quantification for the Armington model, considering estimation error and measurement error simultaneously.
Figure 4: Uncertainty quantification for the trade elasticity in the Armington model.
3.3.2 Relation to Dingel and Tintelnot (2020)
@@
The most relevant paper in the literature on improving counterfactual calculations in quantitative trade and spatial economics is Dingel and Tintelnot (2020), which studies calibration procedures in granular settings. In these settings, individual idiosyncrasies do not wash out and can cause overfitting and poor performance out-of-sample. To deal with this, Dingel and Tintelnot (2020) proposes to, instead of the observed data, either use fitted values obtained using a low-dimensional model or smooth the data using matrix approximation techniques. These procedures can readily be incorporated into Algorithm 1.
Algorithm 2 Uncertainty quantification about $\gamma=g\left(D,\theta\right)$ using Dingel and Tintelnot (2020)
  1. 1.
    Input: noisy data $\tilde{D}$, number of bootstrap draws $B$, coverage level $1-\alpha$ (choose $B$ and $\alpha$ such that $\alpha/2\cdot B\in\mathbb{N}$).
  2. 2.
    For $b=1,...,B$,
    1. (a)
      Sample $D_{b}\sim\pi^{ME}\left(D|\tilde{D}\right)$.
    2. (b)
      Sample $\theta_{b}\sim\pi^{EE}\left(\theta|D_{b}\right)$.
    3. (c)
      Compute $D_{b}^{DT}=s^{DT}\left(D_{b}\right)$.
      1. i.
        For example using a low-dimensional model: $D_{b}^{DT}=\tilde{\mathbb{E}}\left[D_{b}|X\right]$.
      2. ii.
        For example using matrix approximation: $D_{b}=\boldsymbol{U}\boldsymbol{\Sigma}\boldsymbol{V}^{T}\Rightarrow D_{b}^{DT}=\boldsymbol{U}\boldsymbol{\Sigma}_{\tau}\boldsymbol{V}^{T}$, where $\boldsymbol{\Sigma}_{\tau}$ only keeps the first $\tau$ singular values to non-zero.
    4. (d)
@@
      Compute $\gamma_{b}^{DT}=g\left(D_{b}^{DT},\theta_{b}\right).$
  3. 3.
    Sort $\left\{\gamma_{b}^{DT}\right\}_{b=1}^{B}$ to obtain $\left\{\gamma^{DT,(b)}\right\}_{b=1}^{B}$ with $\gamma^{DT,(1)}\leq\gamma^{DT,(2)}\leq...\leq\gamma^{DT,(B)}$.
  4. 4.
    Report $\left[\gamma^{DT,(\alpha/2\cdot B)},\gamma^{DT,(\left(1-\alpha/2\right)\cdot B)}\right].$
The resulting values $\left\{\gamma\right\}_{b=1}^{B}$ can then be interpreted as draws from the posterior of the proposed counterfactual estimator from Dingel and Tintelnot (2020) applied to the data without measurement error.
4 Widely Applicable Default Empirical Bayes Approach

### v5

@@ 3.2 Empirical Bayes and Posterior Distribution
Then, given the estimated hyperparameters $\tilde{\vartheta}$, we can use Bayes’ rule to find the estimated posterior distribution of the true data given the noisy data,
$\pi^{\mathrm{post}}\left(D|\tilde{D},\tilde{\vartheta}\right)=\frac{\pi^{\mathrm{me}}\left(\tilde{D}|D,\tilde{\vartheta}\right)\pi^{\mathrm{prior}}\left(D|\tilde{\vartheta}\right)}{\int\pi^{\mathrm{me}}\left(\tilde{D}|D,\tilde{\vartheta}\right)\pi^{\mathrm{prior}}\left(D|\tilde{\vartheta}\right)dD}.$ (3)
Using this estimated posterior, we can generate draws for the true data given the noisy data. [^3] Note that the measurement error distribution does not have to be mean zero, so also allows for measurement error bias. Nevertheless, even mean zero measurement error can result in bias in the counterfactual prediction of interest. This is automatically taken into account by the Bayesian approach when quantifying uncertainty. Furthermore, the individual measurement error distributions can be arbitrarily correlated in this general setup.
The Bayesian approach allows researchers to incorporate economic knowledge through the prior. For example when considering measurement error in non-negative flows between locations, one can fit a prior centered on a gravity model, which I will do in Section 4.
3.3 Quantifying Uncertainty about $\gamma$
@@
Towards uncertainty quantification, we want to sample from this posterior and report the relevant quantiles. [^4] Note that counterfactual predictions are typically derived as functions of the full system of counterfactual equilibrium variables. Thus, whether the researcher is ultimately interested in a scalar outcome, a relative comparison, or a global average, the mechanics of uncertainty quantification—drawing from the posterior over the true data and solving for equilibrium—remain the same. The entire procedure is summarized in Algorithm 1.
@@
    Input: prior $\pi^{\mathrm{prior}}\left(D|\vartheta\right)$, measurement error model $\pi^{\mathrm{me}}\left(\tilde{D}|D,\vartheta\right)$, noisy data $\tilde{D}$, structural parameter $\theta$, number of bootstrap draws $B$, coverage level $1-\alpha$ (choose $B$ and $\alpha$ such that $\alpha/2\cdot B\in\mathbb{N}$).
@@
    Empirical Bayes estimation step: $\tilde{\vartheta}=\underset{\vartheta}{\arg\max}\int\pi^{\mathrm{me}}\left(\tilde{D}|D,\vartheta\right)\pi^{\mathrm{prior}}\left(D|\vartheta\right)dD.$
  3. 3.
[…]
  4. 4.
    For b=1,,Bb=1,...,B,
@@
      Compute $\gamma_{b}=g\left(D_{b},\theta\right).$
  5. 5.
@@
A natural accompanying point estimator for the procedure in Algorithm 1 is the posterior median, which is always guaranteed to be contained in the reported interval. By contrast, the standard point estimator $g\left(\tilde{D},\theta\right)$, which does not account for measurement error, may lay outside the interval. The posterior median corresponds to an optimal estimate under the estimated posterior and under absolute value loss from a decision-theoretic perspective (see for example Proposition 2.5.5 in robert2007bayesian).
Note that the posterior median does not guarantee bias correction in the frequentist sense. It answers the question: What does a Bayesian believe the counterfactual prediction would have been in the absence of measurement error? While this is a natural and intuitive question to ask, the answer necessarily depends on the prior. But when the prior reflects well-established economic relationships—such as gravity patterns in quantitative trade and spatial models—the posterior median provides a principled estimate of the counterfactual prediction.
3.4 Relation to the Literature
3.4.1 Relation to Measurement Error Literature
The literature on measurement error in nonlinear models is extensive, as reviewed in hu2015microeconomic and schennach2016recent, and the most closely related strand of measurement error literature is that on nonseparable error models (matzkin2003nonparametric; chesher2003identification; hoderlein2007identification; matzkin2008identification; hu2008instrumental; schennach2012local; song2015estimating). However, these results do not apply to my setting.
The key distinguishing feature of the setting in this paper is that the object of interest $\gamma$ directly depends on the correctly measured data, because the equality in Assumption 1 is an exact statement. This is convenient for answering counterfactual questions, and arises because counterfactual questions are typically posed relative to an observed factual situation. In contrast, in standard econometric methods of measurement error, the object of interest is a function of the correctly measured distribution of the data, $\mathcal{P}_{D}$, rather than the actual realized observations, $D$. This leads to the key distinction in Equation (2).
This difference is important because in my setting, it would not suffice to be able to perfectly estimate the distribution $\mathcal{P}_{D}$. For example in a quantitative model of international trade, to answer counterfactual questions we need the realized trade flows, rather than the trade flow distribution from which they are drawn. In contrast, in standard econometric models of measurement error, knowing this distribution would suffice, because the estimands are functionals of the correctly measured distribution of the data. By virtue of that, we need to account for uncertainty about the observations themselves rather than their distribution.
3.4.2 Relation to DingelTintelnot:2025
@@
The most relevant paper in the literature on improving counterfactual calculations in quantitative trade and spatial economics is DingelTintelnot:2025, which studies calibration procedures in granular settings. In these settings, individual idiosyncrasies do not wash out and can cause overfitting and poor performance out-of-sample. To deal with this, DingelTintelnot:2025 proposes to, instead of the observed data, either use fitted values obtained using a low-dimensional model or smooth the data using matrix approximation techniques. Both of these approaches can be cast as special cases of the proposed procedure in Algorithm 1, by choosing a specific prior.
Specifically, the main recommendation is to use a low-dimensional model and is called the covariates-based approach. DingelTintelnot:2025 considers a quantitative spatial model with $L$ individuals. Let $\ell_{ij}$ denote the share of people residing in location $i$ and working in location $j$. The covariates-based approach then interprets the observed commuting shares $\left\{\tilde{\ell}_{ij}\right\}$ as a finite sample from a continuum model. This results in the maximum likelihood model
$\left\{\tilde{\ell}_{ij}\cdot L\right\}|\vartheta\sim\mathrm{Multinomial}\left(\left\{h_{ij}\left(\vartheta\right)\right\}\right),$ (4)
for $\vartheta$ a set of hyperparameters and $h_{ij}\left(\vartheta\right)$ a model function which I discuss further in Appendix B.
[…]
$\tilde{\gamma}^{DT}=g\left(\left\{h_{ij}\left(\tilde{\vartheta}\right)\right\},\theta\right)$
instead of $g\left(\left\{\tilde{\ell}_{ij}\right\},\theta\right)$.
To see how the covariates-based approach from DingelTintelnot:2025 is nested in my Bayesian framework, consider the following prior and measurement error model,
$\begin{cases}\begin{array}[]{c}\mathrm{prior:}\\ \mathrm{measurement\>error:}\end{array}&\begin{array}[]{c}\ell_{ij}|\vartheta\sim\delta_{h_{ij}\left(\vartheta\right)},\quad i,j=1,...,n\\ \left\{\tilde{\ell}_{ij}\cdot L\right\}|\vartheta\sim\mathrm{Multinomial}\left(\left\{\ell_{ij}\right\}\right)\end{array},\end{cases}$ (5)
where $\delta_{h_{ij}\left(\vartheta\right)}$ denotes the Dirac mass at $h_{ij}\left(\vartheta\right)$, implying a degenerate prior. The empirical Bayes step then combines the prior and measurement error model to find
$\left\{\tilde{\ell}_{ij}\cdot L\right\}|\vartheta\sim\mathrm{Multinomial}\left(\left\{h_{ij}\left(\vartheta\right)\right\}\right),$
which overlaps with the model in Equation (4), and uses maximum likelihood estimation to estimate $\vartheta$ by $\tilde{\vartheta}$. This yields the estimated prior and measurement error model
$\begin{cases}\begin{array}[]{c}\mathrm{prior:}\\ \mathrm{measurement\>error:}\end{array}&\begin{array}[]{c}\ell_{ij}|\tilde{\vartheta}\sim\delta_{h_{ij}\left(\tilde{\vartheta}\right)},\quad i,j=1,...,n\\ \left\{\tilde{\ell}_{ij}\cdot L\right\}|\tilde{\vartheta}\sim\mathrm{Multinomial}\left(\left\{\ell_{ij}\right\}\right)\end{array}.\end{cases}$
Using Bayes’ rule we can then find the estimated posterior for the true commuting shares,
$\ell_{ij}|\left\{\tilde{\ell}_{ij}\right\},\tilde{\vartheta}\sim\delta_{h_{ij}\left(\tilde{\vartheta}\right)},\quad i,j=1,...,n.$ (6)
Note that after the empirical Bayes estimation step, since the estimated prior is degenerate, no information is taken from the estimated measurement error model.
The estimated posterior distributions in Equation (6) translate to an estimated posterior for $\gamma$,
@@
$\pi^{\mathrm{post}}\left(\gamma|\left\{\tilde{\ell}_{ij}\right\},\tilde{\vartheta}\right)=\delta_{g\left(\left\{h_{ij}\left(\tilde{\vartheta}\right)\right\},\theta\right)}.$
[…]
Note that if we follow Algorithm 1 exactly, then in step 4a—where we draw from the posterior distributions in Equation (6)—we will obtain the same values in each bootstrap iteration. As a result, the interval constructed in step 6 will collapse to a single point. This outcome is unsurprising, as the procedure in DingelTintelnot:2025 focuses on point estimation and correcting overfitting bias, rather than on quantifying estimation uncertainty. Accordingly, if the sole concern is overfitting bias, their method provides an appropriate approach.
The second recommendation in DingelTintelnot:2025 is to replace the observed data with a smoothed version using matrix approximation techniques. I discuss this approach in Appendix B.
3.5 Estimation Error
The counterfactual prediction of interest will typically depend on a structural parameter $\theta$. It is common in applied work to plug in a fixed value for the structural parameter taken from the literature or obtained through data-driven methods, thus ignoring the uncertainty associated with the estimation process. An exception is adao2017nonparametric, which reports confidence sets for the counterfactual predictions of interest that account for estimation error.
Towards accounting for estimation error for quantitative trade and spatial models in the presence of measurement error, let $\tilde{\theta}$ denote the estimator of the estimand $\theta$. This estimand is usually a function of the distribution of the data $\mathcal{P}_{D}$. This implies that, to address measurement error affecting the structural parameter, one can apply the frequentist measurement error techniques discussed in Section 3.4.1 to find a bias-corrected estimate, though the resulting correction will not admit a Bayesian interpretation.
Alternatively, in Appendix C I outline a fully Bayesian approach that also considers estimation error. Specifically, I assume that the posterior distribution of the structural parameter $\theta$ given the true data $D$ is approximately normal, which is justified under regularity conditions that are closely related to those required for frequentist asymptotic normality. We then have two different posteriors,
[…]
As in Section 3.3, a natural point estimator for the structural parameter is the median of the estimated posterior given the noisy data,
$\pi^{\mathrm{post}}\left(\theta|\tilde{D},\tilde{\vartheta}\right)=\int\pi^{\mathrm{post,ee}}\left(\theta|D\right)\pi^{\mathrm{post,me}}\left(D|\tilde{D},\tilde{\vartheta}\right)dD.$
Using Assumption 1, we can also find the posterior $\pi^{\mathrm{post,ee}}\left(\gamma|D\right)$, and it follows that a natural point estimator for the counterfactual prediction is the median of the estimated posterior
[…]
However, for models satisfying Assumption 1, I am not aware of a frequentist framework that jointly incorporates estimation error and measurement error within a unified procedure that permits both point estimation and uncertainty quantification. By contrast, the proposed Bayesian approach accommodates both sources of uncertainty within a single coherent framework. Moreover, it nests the case in which only measurement error is present: as estimation error vanishes, the procedure naturally reduces to the framework that accounts solely for measurement error. I therefore recommend the Bayesian approach when both sources of uncertainty are relevant, and either the Bayesian or frequentist approach when only estimation error is of concern. In the latter case, the proposed framework can still serve as a robustness check with respect to measurement error.
4 Widely Applicable Default Approach